Alfonzo' pizzeria bakes olive pieces in the outer crust of its 20 inch (diameter) pizza . there is at least one olive piece per inch of crust . how many olive pieces will you get in one slice of pizza ? assume the pizza is cut into eight slices
Answer by nerdybill(7384)

Alfonzo' pizzeria bakes olive pieces in the outer crust of its 20 inch (diameter) pizza . there is at least one olive piece per inch of crust . how many olive pieces will you get in one slice of pizza ? assume the pizza is cut into eight slices
.
Area of pizza = (pi)r^2 = (3.14)10^2 = (3.14)(100) = 314 sq inches
.
Area of a slice of pizza: 314/8 = 39.25 square inches
.
Since there is one olive per square inch:
39.25 olives per slice